A proposed cost-saving device has an installed cost of $815,000. The device will be used in a five-year project but is classified as three-year MACRS property for tax purposes. The required initial net working capital investment is $87,000, the marginal tax rate is 21 percent, and the project discount rate is 11 percent. The device has an estimated Year 5 salvage value of $133,000. What level of pretax cost savings do we require for this project to be profitable? MACRS schedule (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
1) Initial cost of the project = 815000+87000 = $ 902,000.00
2) Terminal non operating cash flow = 133000*(1-0.21)+87000 = $ 192,070.00
PV of terminal non operating cash flows = 192070/1.11^5 = $ 113,984.20
3) PV of depreciation tax shields:
Year Depreciation Rate Depreciation Tax shield on Depreciation PVIF at 11% PV
1 33.33 271639.50 57044.30 0.90090 51391.26
2 44.45 362267.50 76076.18 0.81162 61745.13
3 14.81 120701.50 25347.32 0.73119 18533.74
4 7.41 60391.50 12682.22 0.65873 8354.17
5 0 0.00 0.00 0.59345 0.00
3.69590 140024.29
For NPV = 0,
PV of depreciation tax shields+PV of after tax cost savings+PV of non operating
terminal cash flow-Initial cost = 0
Rearranging the above equation and giving available values we have
PV of after tax cost savings = 902000-113984.20-140024.29 = $ 647,991.51
After tax annual cost savings = 647991.51/3.69590 = $ 175,327.12
Pretax annual cost savings = 175327.12/0.79 = $ 221,933.07 Answer
Answer:
The pretax cost savings should be more than $221,933.07